OpenStudy (anonymous):

What are the next three terms in the sequence -3,6,15,24

OpenStudy (anonymous):

33, 42, 51

OpenStudy (phi):

I would find the difference between each number 6 - -3= 6+3= 9 15-6= 9 24-15= 9 it looks like each term is 9 more than the previous one so the next term past 24 is 9 more than 24 24+9= next number then add 9 to that number to get the next, and so on...